# FedCLEAN: byzantine defense by CLustering Errors of Activation maps in Non-IID federated learning environments

### v1

@@ 5.1.2 Non-IID Data distribution
We distribute the samples of MNIST and FashionMNIST datasets among clients to synthesize a Non-IID data distribution. As mentioned above, we craft distributions that achieve label skew and quantity skew according two scenarios: like in [31] and [54] based on a latent Dirichlet distributions; and following our own custom Non-IID distribution where clients are only allowed to hold samples from two classes. More in details, these two secarios are such as:
  • Dirichlet distribution - we synthesize a population of $C$ clients by drawing a probability vector $\mathbf{q_{n}}\sim D_{C}(\alpha)$ for each class $n\in[1,N]$. Each client is allocated a portion $q_{n,c}$ of the total samples in class $n$. We set $\alpha=1$ [31].
  • Our custom distribution - we enforce the Non-IID condition originally proposed in [41] where each client only possesses samples from two classes. This is achieved by sampling the full datasets using an inverse law with a scale factor of $\alpha$ and an offset of $\gamma$. For the $c^{th}$ client, we draw $D_{s}(c)=\lfloor\frac{\alpha}{c+r}\rfloor+\gamma$ samples without replacement from two arbitrarily selected classes. $r$ ensures $D_{s}$ doesn’t outnumber class samples for small $c$ values. We set $\alpha=2000$, $\gamma=20$, and $r=2$.
Fig. 3 shows an example of such distributions for the MNIST dataset.
5.1.3 Defense setting
We use a CVAE conditioned on class labels. Its encoder and decoder are two-layer MLPs with a hidden layer size of 100 neurons. The number of CVAE training epochs has been fixed to 20. The KLD annealing weight was set to 0 with an increase of 0.5 after 10 epochs. The server $S$ trained its global model copy for 10 warmup epochs, then saved the activation maps from the next 10 epochs with a trigger set containing 250 samples extracted from the test set. We perform trust propagation with $\lambda=0.3$.
Figure 3: Non-IID distributions of the MNIST training samples among 20 clients, colors represent target classes
@@ 5.2 Considered Byzantine attacks
Four attacks tested and parametrized in the literature were tested [12, 24, 6]:
  • Sign-Flipping attacks: Attackers reverse the signs of their model updates by multiplying them by a factor of $-\xi$. This is akin to reversing gradient descent and leading the model in the direction of errors. Let ${\omega_{i}}$ be the weights of the model update after local training. Attackers send back ${\omega^{\prime}_{i}=-\xi\omega_{i}}$.
@@
  • Scaling attack: malicious clients perform local SGD, then simply send back a scaled version of their updates, i.e., ${\omega^{\prime}_{i}=a\omega}$ with $a>1$
Attacks’ parameters $\xi$, $\sigma$, $a$, and $c$ serve as amplification factors that increase the effect of the attack at the cost of making it more detectable For our tests to be nontrivial, we worked with non-upscaled attacks that send updates close to those of benign clients by fixing $\xi=1$, $\sigma=0.1$, $c=0.01$ and $a=10$. We will see that with no defense, these attacks are successful.
@@ 5.3 Experimental results
Table 1: Defense results with the custom distribution
MNIST NAtk AdNse SmVal SgnFlp Scaled FedCLEAN 93.69 84.92 92.78 92.19 91.91 FedCAM 93.02 92.09 93.97 11.29 92.83 FedCVAE 80.82 77.15 11.32 18.68 50.23 FedGuard 61.35 17.67 10.26 11.32 10.26 FLEDGE 93.30 88.06 92.56 92.04 14.19 NoDefense 93.61 18.34 11.32 11.32 12.00 FedCLEAN - 0.00% 0.00% 0.00% 0.00% FedCAM - 0.00% 0.00% 50.60% 0.00% FedCVAE - 0.00% 50.00% 44.60% 1.42% FedGuard - 26.89% 22.15% 12.84% 17.46% FLEDGE - 0.00% 0.00% 0.00% 70.93% NoDefense - - - - - FedCLEAN 0.08% 0.00% 0.00% 0.00% 0.00% FedCAM 47.96% 0.00% 0.00% 47.12% 0.00% FedCVAE 46.79% 0.00% 43.43% 44.60% 1.67% FedGuard 82.01% 69.28% 75.33% 87.28% 82.51% FLEDGE 4.17% 0.00% 0.00% 0.00% 35.08% NoDefense - - - - - FshMNIST NAtk AdNse SmVal SgnFlp Scaled FedCLEAN 69.99 64.52 66.83 66.92 65.83 FedCAM 70.06 60.47 70.83 10.03 57.17 FedCVAE 65.02 64.93 10.07 17.92 60.30 FedGuard 17.29 20.08 10.04 10.04 10.04 FLEDGE 68.78 67.82 74.39 69.95 10.09 NoDefense 71.14 18.86 10.07 13.81 11.97 FedCLEAN - 0.00% 0.00% 0.00% 0.00% FedCAM - 0.00% 0.00% 51.40% 0.00% FedCVAE - 0.00% 51.11% 52.05% 2.68% FedGuard - 27.43% 11.86% 24.11% 18.81% FLEDGE - 0.00% 0.00% 0.00% 70.36% NoDefense - - - - - FedCLEAN 0.68% 0.00% 0.00% 0.00% 0.00% FedCAM 47.48% 0.00% 0.00% 49.51% 0.00% FedCVAE 45.61% 0.00% 43.01% 45.20% 5.98% FedGuard 95.91% 70.90% 83.49% 73.33% 80.52% FLEDGE 3.38% 0.00% 0.00% 0.00% 27.78% NoDefense - - - - -
Table 2: Defense results with the Dirichlet distribution
MNIST NAtk AdNse SmVal SgnFlp Scaled FedCLEAN 95.52 95.39 95.39 95.42 95.51 FedCAM 95.71 95.59 95.57 10.29 95.44 FedCVAE 94.56 93.86 11.32 10.12 17.05 FedGuard 93.20 16.69 11.32 10.33 13.93 FLEDGE 95.58 95.61 95.82 95.31 8.59 NoDefense 95.76 15.63 11.32 10.33 9.76 FedCLEAN - 0.00% 0.00% 0.00% 0.00% FedCAM - 0.00% 0.00% 53.15% 0.00% FedCVAE - 0.00% 56.64% 60.43% 6.54% FedGuard - 21.37% 43.48% 34.14% 45.23% FLEDGE - 0.00% 0.00% 0.00% 50.81% NoDefense - - - - - FedCLEAN 0.04% 0.00% 0.00% 0.00% 0.00% FedCAM 48.04% 0.00% 0.00% 47.97% 0.00% FedCVAE 42.82% 0.00% 49.30% 46.54% 4.90% FedGuard 55.99% 77.73% 58.64% 66.06% 53.67% FLEDGE 0.04% 0.00% 0.00% 0.00% 39.07% NoDefense - - - - - FshMNIST NAtk AdNse SmVal SgnFlp Scaled FedCLEAN 70.86 70.47 68.99 71.06 71.85 FedCAM 74.37 72.55 71.43 10.03 70.19 FedCVAE 68.08 68.09 10.07 17.89 66.75 FedGuard 60.03 15.92 10.07 10.07 12.13 FLEDGE 71.75 70.71 70.40 70.55 9.06 NoDefense 74.39 17.88 10.07 10.01 11.95 FedCLEAN - 0.00% 0.00% 0.00% 0.00% FedCAM - 0.00% 0.00% 52.39% 0.00% FedCVAE - 0.00% 49.35% 56.55% 2.08% FedGuard - 34.65% 31.53% 29.27% 30.50% FLEDGE - 0.00% 0.00% 0.00% 85.71% NoDefense - - - - - FedCLEAN 0.12% 0.00% 0.00% 0.00% 0.00% FedCAM 48.72% 0.00% 0.00% 48.91% 0.00% FedCVAE 44.29% 1.08% 41.91% 45.92% 4.21% FedGuard 70.00% 66.30% 70.67% 72.55% 70.22% FLEDGE 2.59% 0.00% 0.00% 0.00% 13.44% NoDefense - - - - -
Tables 1 and 2 report the results of testing on different distributions of the MNIST and FashionMNIST datasets. We use three standard metrics to measure the effectiveness of each defense system: the global model’s accuracy on the test dataset (ACC), the attacker detection false negative (FN), which refers to the fraction of attackers that manage to get through the defense system, and benign detection false positive (FP) which represents the fraction of blocked benign clients.
For a byzantine attack, the defense requirement R1 requires FN to be 0% in order for the defense to be considered successful. The defense also needs to not disrupt the training (R2), which means ACC should be as close as possible to the benchmark no defense and no attack accuracy, and FP should be as close as possible to 0%.
As we can observe, FedCLEAN is the only defense system that achieves 0% of FN against all types of Byzantine attacks in all tested scenarios. We also see that when attackers are present, FedCLEAN does not block any benign clients (i.e FP=0%). When attackers are absent, it misclassifies less than 1% of benign clients. Additionally, across 20 tested scenarios, FedCLEAN achieves the highest test accuracy in 10 scenarios and matches benchmark accuracy in all 20. But one should note that in all 10 scenarios where FedCLEAN does not achieve the highest accuracy, it achieves 100% attacker detection and 0% or close misclassification, which suggests that the gap in accuracy is due to the random attribution of samples at the start of each experiment.
@@ 6 Conclusion & future work
We introduced FedCLEAN, an autoencoder-based defense system designed to identify malicious clients in Non-IID federated learning environments from updated models’ activation maps. With two core components: A framework for training CVAEs for an FL defense system leveraging model activation maps normalized with GeoMed, DKL annealing, and a realistic trigger set requirement. And a dynamic one-parameter client selection algorithm that minimizes the exclusion of benign clients.
We demonstrate that, in image classification benchmarks, FedCLEAN is robust against Byzantine attacks under a Non-IID client data distribution and achieves top performance in benign client missclassification tests. Future work will focus on broader machine learning applications, backdoor attacks, as well as improving our method by enhancing CVAE training through adaptive regularization, and implementing a fully non-parametric trust propagation algorithm.
6.0.1 Acknowledgements

### v2

@@ 5.1.2 Non-IID Data distribution
Figure 5: Examples of our Non-IID distributions of the MNIST training samples in a scenario with 20 total clients. Colors represent the image classes present in the data.
To synthesize a Non-IID data distribution among clients, we distribute the samples of MNIST and FashionMNIST datasets to achieve:
  • Label skew based on a latent Dirichlet distributions as in [62] and [63] - we synthesize a population of $C$ clients by drawing a probability vector $\mathbf{q_{n}}\sim D_{C}(\alpha)$ for each class $n\in[1,N]$. Each client is allocated a portion $q_{n,c}$ of the total samples in class $n$. We set $\alpha=1$ [62].
  • Quantity skew based on a custom distribution we propose where clients are only allowed to hold samples from two classes - we enforce the Non-IID condition originally proposed in [8] where each client only possesses samples from two classes. This is achieved by sampling the full datasets using an inverse law with a scale factor of $\alpha$ and an offset of $\gamma$. For the $c^{th}$ client, we draw $D_{s}(c)=\lfloor\frac{\alpha}{c+r}\rfloor+\gamma$ samples without replacement from two arbitrarily selected classes. $r$ ensures $D_{s}$ doesn’t outnumber class samples for small $c$ values. We set $\alpha=2000$, $\gamma=20$, and $r=2$.
Figure 5 shows an example of such distributions for the MNIST dataset.
5.1.3 FL-CLEANER defense settings
We use a CVAE conditioned on class labels. Its encoder and decoder are two-layer MLPs with a hidden layer size of 100 neurons. The number of CVAE training epochs has been fixed to 20. The KLD annealing weight was set to 0 with an increase of 0.5 after 10 epochs. The server $S$ trained its global model copy for 10 warmup epochs, then saved the activation maps from the next 10 epochs with a trigger set containing 250 samples extracted from the test set. We perform trust propagation with $\lambda=0.3$ (see Section 4.3). In all our tests, we use all the layers of the model to extract activation maps.
@@ 5.2 Considered Byzantine attacks
Four attacks inspired from literature were tested [11, 12, 16]:
  • Sign-Flipping attacks: Attackers reverse the signs of their model updates by multiplying them by a factor of $\xi$. This is akin to reversing gradient descent and leading the model in the direction of errors. Let $\omega_{i}$} be the weights of the model update after local training. Attackers send back $\omega^{\prime}_{i}=-\xi\omega_{i}$}.
@@
  • Scaling attack: malicious clients perform local stochastic gradient descent, then simply send back a scaled version of their updates, i.e., $\omega^{\prime}_{i}=a\omega$} with $a>1$
The above parameters $\xi$, $\sigma$, $a$, and $c$ control the force or the amplitude of the corresponding attack striking the compromise in between attack efficiency and detectability. For our tests to be nontrivial in terms of detection, we worked with attacks that send updates close to those of benign clients by fixing $\xi=1$, $\sigma=0.1$, $c=0.01$ and $a=10$. We will see that with no defense, these attacks are sufficient to collapse the global model, and that that they are also able to bypass detection by other defenses in literature (see Section 5.4).
5.3 Considered backdoor attacks
For our experiment, we choose two state-of-the-art FL backdoor attacks, which have proven to be successful in non-IID scenarios: The distributed backdoor attack (DBA) [64] and Neurotoxin [65]. Let us recall that backdoor attacks work by adding a pattern (trigger) to training samples, changing their label to a target class, then training on these samples and sending back the corresponding model to the server. The two backdoors we tested work as follows :
  • Distributed backdoor attacks (DBA) [64] : Malicious clients work together to inject the trigger, by each adding a part of the pattern to their samples. When the resulting models are aggregated by the server, the global model associates the entire trigger with the target class
  • Neurotoxin [65] : When training on the poisoned samples, the attackers only update the smallest $k\%$ components of the gradient, which are less likely to be overwritten during aggregation.
We test these attacks with a $10\times 10$ pixels white square pattern for both MNIST and FashionMNIST, split into four $5\times 5$ squares for DBA. For Neurotoxin, we set $k=95\%$.
@@ 5.4 Experimental results
Figure 6: Defense Results on MNIST considering two non-IID distributions against Byzantine attacks: Additive noise, Same value, Sign Flipping and Scaled attacks of parameters $\xi=1$, $\sigma=0.1$, $c=0.01$ and $a=10$ respectively.
Figure 7: Defense Results on FashionMNIST considering two non-IID distributions against Byzantine attacks: Additive noise, Same value, Sign Flipping and Scaled attacks of parameters $\xi=1$, $\sigma=0.1$, $c=0.01$ and $a=10$ respectively.
For Byzantine attacks, we use three standard metrics to measure the effectiveness of each defense system: the global model’s accuracy on the test dataset (ACC); the attacker detection recall (Recall) which refers to the percentage of attackers successfully blocked by the defense system; and, benign detection false positive (FPR) which represents the fraction of blocked benign clients. Figures 6 and 7 report the results of testing on the two distributions of the MNIST and FashionMNIST datasets. For byzantine attacks, recall that a single attacker is sufficient to collapse the global model in just one FL round. Thus, a Recall value of 100% is required in order the defense to be considered successful and reach the defense requirement R1 (see Section 3.2). The defense also needs to not disrupt the training (R2), which means ACC should be as close as possible to the benchmark no defense and no attack accuracy, and FPR should be as close as possible to 0%. As we can observe in Figures 6 and 7, compared to the most recent effective methods in literature (see Section 1.1), FL-CLEANER is the only defense system that achieves 100% Recall against all types of Byzantine attacks in all tested scenarios. We can also see that when attackers are present, FL-CLEANER does not block any benign clients (i.e FPR=0%). When attackers are absent, it misclassifies less than 1% of benign clients. Additionally, across 20 tested scenarios, FL-CLEANER achieves the highest test accuracy in 10 scenarios and matches benchmark accuracy in all 20. Note that one reason why FL-CLEANER does not reach the highest accuracy achieved in the scenario with no-defense and no attack may come from the fact that in the experiments of FL-CLEANER only 70% of the training data are used for training. The rest is used to simulate malicious clients for test purposes while in the scenario with no-defense and no attack, all data are used for training. Regardless, FL-CLEANER achieves 100% attacker detection and 0% or close miss-classification in each of these cases.
It is important to note that other defenses were implemented according to the parameters provided by their authors. For FedCVAE for example, this meant using the whole test set, i.e, 10 000 images, to train the its CVAE. It can be seen that FL-CLEANER achieves better performance with a much smaller trigger set of 250 samples.
Defense DBA Neurotoxin
ASR ($\downarrow$) ACC ($\uparrow$) ASR ($\downarrow$) ACC ($\uparrow$)
NoDefense 95.4% 21.2% 9.1% 92.8%
FLAME 0.0% 9.8% 0.0% 16.8%
FLDetector 11.9% 85.7% 9.6% 93.9%
FL-CLEANER 0.2% 95.0% 0.0% 95.1%
Table 1: Attack Success Rate (ASR) and Accuracy (ACC) for different defenses and attacks on MNIST. Arrows represent ideal outcome.
Defense DBA Neurotoxin
ASR ($\downarrow$) ACC ($\uparrow$) ASR ($\downarrow$) ACC ($\uparrow$)
NoDefense 6.1% 65.8% 100% 59.6%
FLAME 9.1% 42.1% 23.4% 54.1%
FLDetector 16.3% 75.0% 43.9% 71.4%
FL-CLEANER 0.2% 67.2% 0.3% 69.1%
Table 2: Attack Success Rate (ASR) and Accuracy (ACC) for different defenses and attacks on FashionMNIST. Arrows represent ideal outcome.
Concerning backdoor attacks, let us recall that the goal of the attacker is to achieve high ASR while preserving the model’s performance on the main task (see Section 3.1). Thus, we evaluate our defense on the impact it has on this metric. To evaluate this, we first test the backdoors in the absence of our defense, and track their ASR over the training epochs. Then, we do the same with a defense applied to compare. In addition, we report the final attacker success rate and test accuracy at the end of the run in separate tables. In our testing, attackers struggled to insert the backdoor when training on the more restrictive custom distribution. So we focused our tests on the Dirichlet distribution.
For comparison with other defenses, we decided not to take into account backdoor defenses the strategy of which includes a tolerance mechanism sometimes letting one to several attackers get through the defense [27, 15]. These approaches are by design not byzantine-robust. Indeed, most byzantine attacks only need one attacker to get through. We choose to focus on two backdoor defenses that filter out all detected attackers : FLAME [28] and FLDetector [29]. They are also considered as the state-of-the-art [66]
Figures 9 and 9 show the evolution of backdoor accuracy (ASR) over a 50-round FL run. Tables 1 and 2 report the final attacker success rate (ASR) and test accuracy (ACC) observed at the end of these runs. We see that :
  1. A
    FL-CLEANER cancels the effect of both attacks : After a few rounds, we see that it becomes impossible for the attacker to maintain an ASR not nearing 0%. This validates the backdoor defense requirement R3.
  2. B
    FL-CLEANER is the only effective defense that does not harm the model accuracy : As seen in Tables 1 and 2, when other defenses manage to mitigate the backdoor attack, it comes at the cost of the model’s accuracy. This is due to the defenses blocking too many benign clients. Which makes FL-CLEANER the only defense to validate the defense requirement R2.
Figure 8: Backdoor Attacker Success rates on MNIST
with and without defense (smoothed)
Figure 9: Backdoor Attacker Success rates on Fashion-
MNIST with and without defense (smoothed)
@@ 6 Conclusion & future work
In this paper, we introduced FL-CLEANER the first method robust against Byzantine and backdoor attacks under a Non-IID client data distribution. Its originality stands on an autoencoder-based defense system designed to identify malicious clients in Non-IID federated learning environments from updated models’ activation maps with other core components: a framework for training CVAEs for an FL defense system leveraging model activation maps normalized with GeoMed, DKL annealing, and a realistic trigger set requirement. Another strength of our proposal is that it depends on a dynamic one-parameter client selection algorithm that minimizes the exclusion of benign clients. We demonstrated that, in image classification benchmarks, FL-CLEANER is robust against Byzantine attacks as well as effective to mitigate tested state-of-the-art backdoor attacks, achieving top performance in benign client misclassification tests compared to current defenses from the literature. Future work will focus on broader machine learning applications, as well as improving our method by enhancing CVAE regularization in the training, and implementing an improved trust propagation algorithm.
Acknowledgments